# A glance to Luttinger liquid and its platforms

### v1

@@ VI.2 Spin-charge separation in 1D atomic fermions
Unlike three-dimensional (3D) systems whose low-energy excitations are fermionic quasi-particles, the low-energy excitations of one-dimensional (1D) fermions are collective bosonic spin- and charge-density waves (SDW/CDW) that disperse linearly, as described by the Tomonaga-Luttinger liquid (TLL) theory[110, 156, 56]. A peculiar result is that the SDW and the CDW of an interacting 1D Fermi gas propagate at different speeds, thus showing a spatial separation of spin and charge excitations in the gas. Recently, a series of experiments with ultracold atoms in an optical lattice was performed on a single-site resolved 1D Hubbard chain, leading to the observation of the fractionalization of spin and charge quantum numbers[73], including the modification of the SDW wavevector by density-doping and by spin-polarization[133], and the study of simultaneous spin and charge dynamics after performing a quench [158]. Quite remarkably these studies allowed to see directly the non-local string structure predicted by the theory between the spinons and holons. However, the excitations produced by the quench are not in the linearly dispersing regime, so they can only be qualitatively be compared with the TLL theory.
@@
Figure 6: Spin and charge excitations via Bragg spectroscopy. (A) Relevant transitions and laser detunings for spin ( S, violet) and charge ( C, red) excitations in 6Li atoms. (B) Bragg beams (1, 2) oriented to produce a standing wave along the tube axis. A momentum transfer $\vec{q}=\vec{k_{1}}-\vec{k_{2}}$ is imparted to the atoms. (C,D) Schematic diagram of the charge and spin excitations, showing an excitation of a particle-hole pair, or a spinon pair. At the bottom is shown the effect on the total density $\rho(x)$ and spin density $\sigma$(x).
@@ VII.1 The contribution of nonlinearities
The TL liquid theory has served as a useful paradigm for describing one-dimensional (1D) quantum fluids in the limit of low energies. This theory is based on a linearization of the dispersion relation. In the phenomenological LL approach, the conventional justification for such simplification is irrelevance of the nonlinearity in the renormalization group sense. Progress in understanding 1D quantum fluids beyond the low-energy limit has been proposed [79], and studied experimentally and compared with theory[141, 26]. The irrelevant terms hardly affect the fermion propagator away from the singular lines in space-time, $x\pm vt\gg\sqrt{t/m*}$. The vicinity of these lines defines the singular behavior of the spectral function. It has been shown for all spinless 1D fermionic models with short range interactions that the single-particle spectral function at $p<<k_{F}$ is universal. In the vicinity of the Fermi wave vector $+k_{F}$ and $p,\omega>0$, the spectral function $A(p,\omega)$ is a universal function of a single argument:
@@
where $p$ is measured from the closest Fermi point, and $\hbar=1$. The function $A(\epsilon)$ is very different from the LL theory predictions, yet it depends only on the LL parameter K. The asymptotic behaviour of $A(\epsilon)$ for $\epsilon>>1$ recovers the LL theory predictions, but at $|\epsilon\pm 1|\ll<<1$ the spectral function is described by a power-law with new exponents. The exponents are different from the predictions of the LL theory yet can still be analytically expressed in terms of $K$. The exponent can be determined numerically by the nonlinear dynamics in nonequilibrium Fermi gases[1, 12, 13].
The methodology described above, which combines perturbation theory relative to interaction strength, and innovative techniques for handling finite-size properties of integrable models, is applicable to a broad spectrum of 1D fluids, including 1D spin liquids, electrons in quantum wires, and cold atoms confined in 1D traps. Such an approach additionally demonstrates how nonlinear dispersion leads to finite particle lifetimes and influences the transport properties of 1D systems at finite temperatures.
@@ VII.4 Quantum critical Luttinger liquid behavior and spinon confinement
Low dimensional quantum magnets are interesting systems because of the emerging collective behavior arising from strong quantum fluctuations which can lead to LL behavior. The one-dimensional (1D) S=1/2 Heisenberg antiferromagnet is a paradigmatic example whose excitations, known as spinons[42], carry spin $S=1/2$, half the spin of a magnon. In the Luttinger liquid phase, these fractional modes are deconfined and have linear dispersion[9, 22]. They can be reconfined by the application of a staggered magnetic field or a dimerization[67]. Spinons and their confinement have been observed in a variety of experimental systems[92, 91, 103, 8, 54, 170, 157, 55]. The spinon description of the dynamical structure factor extends beyond the low-energy long-wavelength limit[103].
@@ VIII Outlook and conclusions
Moreover, the field stands to benefit greatly from continued advances in cold atom experiments. The extreme tunability of these platforms enables precise control over interactions, dimensionality, and particle statistics, providing a powerful testbed for exploring quantum-critical behavior, out-of-equilibrium dynamics, and disorder. The ability to simulate interacting Luttinger liquids in optical lattices or confined geometries could deliver fresh insights into the behavior of TLLs under conditions that were previously difficult or impossible to achieve in solid-state systems.
@@ Acknowledgements
R.C. was partly supported by the PNRR MUR project PE0000023 NQSTI (TOPQIN and SPUNTO). I.B., R.C., E.O. and T. G. would like to thank the Institut Henri Poincaré (UAR 839 CNRS Sorbonne Université) and the Lab Ex CARMIN(ANR-10 LABX-59-01)for their support. T. G. was supported in part by the Swiss National Science Foundation under grants 200020-188687 and 200020-219400. B. W. acknowledges the support of the National Research Foundation (NRF) Singapore, under the Competitive Research Program “Towards On-Chip Topological Quantum Devices” (NRF-CRP21-2018-0001), with further support from the Singapore Ministry of Education (MOE) Academic Research Fund Tier 3 grant (MOE-MOET32023-0003) “Quantum Geometric Advantage”. T. D. acknowledges suport from the Centre of Excellence for Engineered Quantum Systems, an Australian Research Council Centre of Excellence, CE110001013. M. K. acknowledges the financial support of the Slovenian Research and Innovation Agency through the programme No. P1-0125 and the project No. J1-2456. RGH was funded in part by the National Science Foundation (US), grant no. 2309362

### v2

@@ VI.2 Spin-charge separation in 1D atomic fermions
Unlike three-dimensional (3D) systems whose low-energy excitations are fermionic quasi-particles, the low-energy excitations of one-dimensional (1D) fermions are collective bosonic spin- and charge-density waves (SDW/CDW) that disperse linearly, as described by the Tomonaga-Luttinger liquid (TLL) theory[118, 167, 61]. A peculiar result is that the SDW and the CDW of an interacting 1D Fermi gas propagate at different speeds, thus showing a spatial separation of spin and charge excitations in the gas. Recently, a series of experiments with ultracold atoms in an optical lattice was performed on a single-site resolved 1D Hubbard chain, leading to the observation of the fractionalization of spin and charge quantum numbers[78], including the modification of the SDW wavevector by density-doping and by spin-polarization[143], and the study of simultaneous spin and charge dynamics after performing a quench [173]. Quite remarkably these studies allowed to see directly the non-local string structure predicted by the theory between the spinons and holons. However, the excitations produced by the quench are not in the linearly dispersing regime, so they can only be qualitatively be compared with the TLL theory[12].
@@
Figure 6: Spin and charge excitations via Bragg spectroscopy. (A) Relevant transitions and laser detunings for spin (S, violet) and charge (C, red) excitations in 6Li atoms. (B) Bragg beams (1, 2) oriented to produce a standing wave along the tube axis. A momentum transfer $\vec{q}=\vec{k_{1}}-\vec{k_{2}}$ is imparted to the atoms. (C,D) Schematic diagram of the charge and spin excitations, showing an excitation of a particle-hole pair, or a spinon pair. At the bottom is shown the effect on the total density $\rho(x)$ and spin density $\sigma(x)$. The figure is reproduced by Ref.[151]
@@ VII.1 The contribution of nonlinearities
The TL liquid theory has served as a useful paradigm for describing one-dimensional (1D) quantum fluids in the limit of low energies. This theory is based on a linearization of the dispersion relation. In the phenomenological LL approach, the conventional justification for such simplification is irrelevance of the nonlinearity in the renormalization group sense. Progress in understanding 1D quantum fluids beyond the low-energy limit has been proposed [84], and studied experimentally and compared with theory[8, 124, 90, 172, 151, 28]. The irrelevant terms hardly affect the fermion propagator away from the singular lines in space-time, $x\pm vt\gg\sqrt{t/m*}$. The vicinity of these lines defines the singular behavior of the spectral function. It has been shown for all spinless 1D fermionic models with short range interactions that the single-particle spectral function at $p<<k_{F}$ is universal. In the vicinity of the Fermi wave vector $+k_{F}$ and $p,\omega>0$, the spectral function $A(p,\omega)$ is a universal function of a single argument:
@@
where $p$ is measured from the closest Fermi point, and $\hbar=1$. The function $A(\epsilon)$ is very different from the LL theory predictions, yet it depends only on the LL parameter K. The asymptotic behaviour of $A(\epsilon)$ for $\epsilon>>1$ recovers the LL theory predictions, but at $|\epsilon\pm 1|\ll 1$ the spectral function is described by a power-law with new exponents[171, 170]. The exponents are different from the predictions of the LL theory yet can still be analytically expressed in terms of $K$. The exponent can be determined numerically by the nonlinear dynamics in nonequilibrium Fermi gases[1, 14, 15].
The methodology described above, which combines perturbation theory relative to interaction strength, and innovative techniques for handling finite-size properties of integrable models, is applicable to a broad spectrum of 1D fluids, including 1D spin liquids, electrons in quantum wires[171, 170], and cold atoms confined in 1D traps. Such an approach additionally demonstrates how nonlinear dispersion leads to finite particle lifetimes and influences the transport properties of 1D systems at finite temperatures.
@@ VII.4 Quantum critical Luttinger liquid behavior and spinon confinement
Low dimensional quantum magnets are interesting systems because of the emerging collective behavior arising from strong quantum fluctuations which can lead to LL behavior. The one-dimensional (1D) S=1/2 Heisenberg antiferromagnet is a paradigmatic example whose excitations, known as spinons[46], carry spin $S=1/2$, half the spin of a magnon. In the Luttinger liquid phase, these fractional modes are deconfined and have linear dispersion[10, 24]. They can be reconfined by the application of a staggered magnetic field or a dimerization[72]. Spinons and their confinement have been observed in a variety of experimental systems[100, 99, 111, 9, 59, 185, 168, 60]. The spinon description of the dynamical structure factor extends beyond the low-energy long-wavelength limit[111]. Spinon and holon excitations are also present in fermion systems[124], and holon contributions are expected to dominate the momentum distribution[169] near $3k_{F}$.
@@ VIII Outlook and conclusions
Moreover, the field stands to benefit greatly from continued advances in cold atom experiments. The extreme tunability of these platforms enables precise control over interactions, dimensionality, and particle statistics, providing a powerful testbed for exploring quantum-critical behavior, out-of-equilibrium dynamics, and disorder. The ability to simulate interacting Luttinger liquids in optical lattices or confined geometries could deliver fresh insights into the behavior of TLLs under conditions that were previously difficult or impossible to achieve in solid-state systems[37].
@@ Acknowledgements
R.C. was partly supported by the PNRR MUR project PE0000023 NQSTI (TOPQIN and SPUNTO). I.B., R.C., E.O. and T. G. would like to thank the Institut Henri Poincaré (UAR 839 CNRS Sorbonne Université) and the Lab Ex CARMIN(ANR-10 LABX-59-01)for their support. T. G. was supported in part by the Swiss National Science Foundation under grants 200020-188687 and 200020-219400. B. W. acknowledges the support of the National Research Foundation (NRF) Singapore, under the Competitive Research Program “Towards On-Chip Topological Quantum Devices” (NRF-CRP21-2018-0001), with further support from the Singapore Ministry of Education (MOE) Academic Research Fund Tier 3 grant (MOE-MOET32023-0003) “Quantum Geometric Advantage”. T. D. acknowledges support from the Centre of Excellence for Engineered Quantum Systems, an Australian Research Council Centre of Excellence, CE110001013. M. K. acknowledges the financial support of the Slovenian Research and Innovation Agency through the programme No. P1-0125 and the project No. J1-2456. RGH was funded in part by the National Science Foundation (US), grant no. 2309362
Author contributions
The authors contributed equally to all aspects of the article.
Competing interests
The authors declare that they have no competing interests.